The perimeter of the rectangle is 118 feet. Find the length of each side. the length of one side is 10x-5 and the width is 6x.
Answer by mananth(16946)

the length of one side is 10x-5 and the width is 6x.
Perimeter = 2(L+W)
118=2(10x-5+6x)
118=2(16x-5)
/2
59=16x-5
59+5=16x
64=16x
/16
x=4
one side = 10x-5 = 40-5 = 35feet
other side = 6x =24 feet